OpenStudy (anonymous):

How do you change 76 degrees 22' 45'' into a fraction (Keep as a fraction)

jimthompson5910 (jim_thompson5910):

22' means 22 minutes 45" means 45 seconds

jimthompson5910 (jim_thompson5910):

60 minutes = 1 degree 1 minute = 1/60 degree 22 minutes = 22/60 degree ------------------------------------------------------- 60 seconds = 1 minute 60*60 seconds = 60*1 minute 3600 seconds = 60 minutes 3600 seconds = 1 degree 1 second = 1/3600 degree 45*1 second = 45*1/3600 degree 45 seconds = 45/3600 degree

jimthompson5910 (jim_thompson5910):

so this all means that 76 degrees 22' 45'' turns into 76 degrees + 22/60 degree + 45/3600 degree I'll let you finish up